# The frame-dragging vector potential on galaxy scales from DM-only Newtonian $N$-body simulations

### v2

@@ Acknowledgements
We wish to thank Mario G. Lattanzi for useful comments and his continuous support. We also thank James Trayford, Christopher Lovell, Daniel B. Thomas, and Kazuya Koyama for helpful discussions. This work has made use of the SCIAMA supercomputer of the University of Portsmouth’s Institute of Cosmology and Gravitation.
@@
Data underlying this article will be shared on reasonable request to the corresponding author.
@@ A.1 Dependence on the DTFE grid resolution
Here, the power spectra of the source terms of the gravito-magnetic potential are checked to have achieved convergence while varying the grid resolution of the DTFE. Figure 20 shows that convergence is guaranteed for $N_{\mathrm{grid}}>512$ for all quantities.
Figure 20: Power spectra of the velocity and momentum gradients, as detailed in Figure 3 for the simulation TNG300-3-Dark at $z=0$, for different grid resolutions of the DTFE: $N_{\mathrm{grid}}=128$ (cyan), $256$ (purple), $384$ (orange), $512$ (magenta), $768$ (green), $896$ (red), and $1024$ (blue). The divergence and vorticity of the velocity are represented as solid and dashed lines, respectively, while the divergence and vorticity of the momentum are shown with dotted and dash-dotted lines. The linear matter power spectrum is plotted as a reference (grey solid line), together with the non-linear matter power spectra measured from the simulation for the corresponding $N_{\mathrm{grid}}$.

### v3

@@ Acknowledgements
We express our gratitude to the Reviewer for their time and insightful comments that contributed to improving the original manuscript. We wish to thank Mario G. Lattanzi for useful comments and his continuous support. We also thank James Trayford, Christopher Lovell, Daniel B. Thomas, and Kazuya Koyama for helpful discussions. This work has made use of the SCIAMA supercomputer of the University of Portsmouth’s Institute of Cosmology and Gravitation. For the purpose of open access, we have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.
@@
Supporting research data are available on reasonable request from the authors.
@@ A.1 Dependence on the DTFE grid resolution
Here, the power spectra of the source terms of the gravito-magnetic potential are checked to have achieved convergence while varying the grid resolution of the DTFE. Figure 20 shows that convergence is guaranteed for $N_{\mathrm{grid}}=1024$ for all quantities, as the green ($N=768$), red ($N=896$), and blue ($N=1024$) lines are very close to each other.
Figure 20: Power spectra of the velocity and momentum gradients, as detailed in Figure 3 for the simulation TNG300-3-Dark at $z=0$, for different grid resolutions of the DTFE: $N_{\mathrm{grid}}=128$ (cyan), $256$ (purple), $384$ (orange), $512$ (magenta), $768$ (green), $896$ (red), and $1024$ (blue). The divergence and vorticity of the velocity are represented as solid and dashed lines, respectively, while the divergence and vorticity of the momentum are shown with dotted and dash-dotted lines. The linear matter power spectrum is plotted as a reference (grey solid line), together with the non-linear matter power spectra (dot-dot-dashed line) measured from the simulation for the corresponding $N_{\mathrm{grid}}$.